Find the equation of y, given the derivative and the indicated point on the curve. oy = 2(x - 1)
Answer #1

Solhi o = 2(x-1) dy = 2(x-1)da Integrating both sides, we have l 2(2-1) dai tc., c be the a constant of integra- tion : > y =